OpenStudy (imnotgoodatmath):

What is the y-intercept of the line given by the equation y equals two x minus nine? The y-intercept is 0. The y-intercept is Negative four point five. The y-intercept is Negative nine. The y-intercept is 18.

OpenStudy (imnotgoodatmath):

@mathstudent55

OpenStudy (anonymous):

y = 2x-9 The Y-intercept is wherever the graph reaches x=0. Set x to 0 and plug it in: y =2(0)-9 y= -9

OpenStudy (anonymous):

And y=mx+b is already y-intercept form, where b is the y-intercept and m is the slope